I, tI 2-Ig=0 2Izt Iz=-7 4I t Iz=21

I got these three final equations of kirchoffs law but need help solving for the current value for each I.
Answer #1

I+ Iz - Iz = 0 - I = I3-Iq-11) 21, + Iz :.7 - (2) :4 I, + Iz :21 - (3) By Dutting value of I, in (?) -) 4.(Ix-I.) +I3=21 ,1.